OpenStudy (anonymous):

y+12=5y-4. What do i do? the direction is "sole each eqation or inequality. Each solution is an integer"

OpenStudy (anonymous):

First you have to subtract y from both sides. y + 12 = 5y - 4 -y -y ----------------- 12 = 4y - 4 Now you have to add 4 to both sides. +4 +4 ----------- 16 = 4y Then divide. /4 /4 --------- y = 4 Hope this helped!

OpenStudy (anonymous):

Thankyou. i didnt know it was that simple. i thought it was something complacated

OpenStudy (anonymous):

lol I know that's how they always try to put it